OpenStudy (anonymous):

If 3a - 9 = 6, what is the value of 4a - 1?

OpenStudy (anonymous):

show all steps

OpenStudy (anonymous):

add 9 to get 3a=15 divide by 3 to get a=5 plug a=5 into the second equation to get 4(5)-17=20-17=3